OpenStudy (anonymous):

MilkPoppers 3. a. What is the common ratio of the sequence 2, -6, 18, -54,.... b. What are the next three terms in the sequence?

OpenStudy (the_fizicx99):

-54/18 = ?

OpenStudy (anonymous):

-3 , right?

OpenStudy (the_fizicx99):

Find the rate first, -54 / 18 = ? 18 / -6 = ? and so on, Then plug in what you know into a formula or simply enter it into a calculator. Which one do you want to go with?

OpenStudy (the_fizicx99):

The explicit formula is: f(n) = f(1) • r^n-1 Use alt 7 to use the " • " dot :P

OpenStudy (anonymous):

Okay, I don't understand.. I'm lost.. lol. Okay, so -54 / 18 = -3

OpenStudy (the_fizicx99):

Yes, that's your rate.

OpenStudy (anonymous):

Ohh, okay. So how do I find out that the next three numbers?

OpenStudy (the_fizicx99):

f(n) = f(1) • r^n-1 find the 5th 6th and 7th term I'll do the 5th for you, f(5) = f(1) • -3^5-1 f(5) = 2 • -3^4 f(5) = 2 •81 f(5) = 162

OpenStudy (anonymous):

Okay, let me see here. It might take a moment. lol

OpenStudy (the_fizicx99):

Geddit?

OpenStudy (the_fizicx99):

No?

OpenStudy (anonymous):

f(5) = f(1) -3^6-1 f(5) = 2 -3^5 f(5) = 2 -243 f(5) = -486 ?

OpenStudy (the_fizicx99):

Not f(5).. it's f(6) f(6) = -486 Yes.

OpenStudy (anonymous):

Ohh, I got it now. Thank you so much. Algebra isn't my thing anymore. lol

OpenStudy (the_fizicx99):

:p np

OpenStudy (anonymous):

F(7) = f(1) -3^7 – 1 f(7) = 2 -3^6 f(7) = 2 729 F(7) = 1458

OpenStudy (the_fizicx99):

Yup

OpenStudy (anonymous):

Thank you, I get it now. :D

OpenStudy (the_fizicx99):

yw

OpenStudy (the_fizicx99):

Gl :***

OpenStudy (anonymous):

In those kind of problems do you always do the minus 1 in the first part of the equation?

OpenStudy (the_fizicx99):

That's the explicit formula you can always do the recursive. f(n) = f(n – 1) • r

OpenStudy (anonymous):

Ohh, okay. lol thanks.

OpenStudy (the_fizicx99):

k